OpenStudy (anonymous):

please solve x^3-8=0

OpenStudy (usukidoll):

add 8 to both sides....

OpenStudy (usukidoll):

and then multiply by 1/3 exponential 1/3

OpenStudy (usukidoll):

then rewrite the right side in radical form and take the cube root of 8

OpenStudy (usukidoll):

3 should be written outside the sq root and since it's just 1 you don't have to write 8 to the first power

OpenStudy (anonymous):

and ans is 2???

OpenStudy (usukidoll):

yes plug it back into the original equation.

hartnn (hartnn):

you know there will be 3 roots of x ? you only need integer one ?

OpenStudy (usukidoll):

I need to plug my pc back in brb

OpenStudy (anonymous):

why can't we use formula a^3-b^3=(a-b)(a^2+ab+b^2)????

hartnn (hartnn):

you should use that formula...

OpenStudy (anonymous):

if we use that formula we will get one real and two complex roots

hartnn (hartnn):

correct

OpenStudy (anonymous):

thats what i want to ask.... x^3=8 x^3=2^3 x=2 give only real root but applying formula give 1 real and 2 complex roots why this difference come on applying two different(but correct) methods??????

hartnn (hartnn):

"x^3=2^3 x=2 " this is correct but incomplete. x^3=2^2 x^3-2^3=0 now apply a^3 - b^3 formula this method is correct and complete.

OpenStudy (anonymous):

\[x^3 -8 \\ x=\sqrt[3]{8}\\ x=\sqrt[3]{2^3}\\ x = 2\]

OpenStudy (austinl):

I shall join the long line of people answering. \(\large{x^3-8=0}\) \(\large{x^3=8}\) \(\large{x=\sqrt[3]{8}}\) \(\large{x=2}\) I feel happy now.

OpenStudy (anonymous):

Go for hartnn's explanation, he did the complexes as well

OpenStudy (zzr0ck3r):

lol there are 3 solutions! listen to @hartnn

OpenStudy (zzr0ck3r):

everyone else stop saying 2 is the answer...

OpenStudy (anonymous):

what is this, a "who draws better latex" competition ?!

OpenStudy (anonymous):

Let p(x)=x^3-8 Since you've known x=2 is a root Then (x-2) is a factor of p(x) Do a synthetic division  Then use the quotient and use any factoring methods to get your complex roots

OpenStudy (austinl):

\(\Huge{\color{red}{\text{Y}}}\Large{\color{blue}{\text{E}}}\LARGE{\color{green}{\text{S}}}\) @Luis_Rivera

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Do you understand rosy?

OpenStudy (zzr0ck3r):

\[a^3-b^3=(a-b)(a^2+ab+b^2)\\x^3-2^3=0\\(x-2)(x^2+2x+4)=0\]now solve these\[x=2\\and\\(x+1)^2=-3\\x=\pm(i\sqrt{3}-1)\]

OpenStudy (zzr0ck3r):

ok back to vacation in paris:) see ya suckers

OpenStudy (anonymous):

Hey mind your language zzr0ck3r :)

OpenStudy (anonymous):

by the way, @Z3R0, i wonder what problem you're working on ? cause it certanly aint this one !

OpenStudy (anonymous):

These are the basic steps for beginners like him

OpenStudy (nincompoop):

diving in! (x-2) (x^2+2 x+4) = 0